Allison:

Ben bowled 124 and 219 in his first two games. What must he bowl in his third game to have an average of at least 180 ​?

QuestionCoveBot:

So, we have to add the two numbers together then divide by 2. So 219 + 124 would be?

Darky2:

Yeah i think thats it ._.

Vocaloid:

There are *three* scores so you cannot simply take the average of two numbers. Let the third score be x. The average of all three scores is the sum of the three scores, divided by 3, or in other words, (124 + 219 + x)/3 and let this be greater than/equal to 180.

Allison:

197

Vocaloid:

yup good

TutorJohnfoster:

let the third unknown number be x Average=\[\frac{ The\space total\space value\space of \space the \space items }{number\space of \space item ? }\] number of item=3 The total value of the item=124+219+x Average=180 Thus \[180=\frac{ 124+219+x }{ 3}\] 180*3=124+219+x 540=343+x 540=343+x 540-343=x 197=x The third number is 197